# Sequential Change Detection under Markov Setup with Unknown Prechange and Postchange Distributions

### v1

@@ II-A Previous Work
Here, $\lambda$ was first used by the authors in order to ensure a negative drift in the test statistic in the pre-change regime, as is the case in the classical Page’s test in the parametric setup.
@@ III-A Estimating Pre-change Distribution
$\hat{\mu}_{0}(x_{1}^{n_{0}})=\frac{\hat{\mu}_{0}(x_{1},x_{2})\hat{\mu}_{0}(x_{2},x_{3})\dots\hat{\mu}_{0}(x_{n_{0}-1},x_{n_{0}})}{\hat{\mu}_{0}(x_{2})\dots\hat{\mu}_{0}(x_{n_{0}})}$
@@
Here, $N(b)\mid x_{1}^{n_{0}}$ represents the number of times the symbol $b$ appears in the sequence $x_{1}^{n_{0}}$, whereas $N(b,a)\mid x_{1}^{n_{0}}$ represents the number of times the symbols $b,a$ jointly appear in the sequence $x_{1}^{n_{0}}$.Also, $\hat{\mu}_{0}(b)$ is the fraction of times $b$ appears in the sequence $x_{1}^{n_{0}}$, whereas $\hat{\mu}_{0}(b,a)$ is the fraction of times $b,a$ appear in the sequence $x_{1}^{n_{0}}$ jointly.
We assume that $n_{0}$ is large enough so that each symbol $b\in\mathcal{X}$ which has non-zero probability under ${\mu}_{0}$ has appeared at least once in $x_{1}^{n_{0}}$ and symbol pair $(b,a)\in\mathcal{X}^{2}$ which have non-zero probability under ${\mu}_{0}$ have appeared at least once in $x_{1}^{n_{0}}$ (else $\hat{\mu}_{0}$ is not an accurate estimate of ${\mu}_{0}$).
Since the process is stationary,irreducible and aperiodic Markov Chain satisfying Berk’s Mixing Conditions,by Ergodic Theorem we can say that there exists an $n_{0}$ such that,
@@ III-C Behaviour Under Prechange Distribution
Lemma 2: If $x_{1}^{n_{0}+n}\in E_{n}$, then $f_{n}(x_{n_{0}+1}^{n_{0}+n})$ is bounded above for all $n$ as $f_{n}\leq\log\beta$ where $\beta=\beta^{\prime}\cdot\beta^{\prime\prime},\beta^{\prime}=\frac{\mu_{0}(a\mid b)}{\mu_{0}(a\mid b)-\delta},\beta"=\frac{\mu_{0}(b)}{\mu_{0}(b)-\delta^{\prime}}$ and $b,a$ is such that $b,a$ is the minimum probability symbol pair and $b$ is such that $b$ is the minimum probability symbol.As $\delta\to 0$$\delta^{\prime}\to 0$ and $n_{0}\to\infty$, we have $\beta\to 1$.
@@
The terms on RHS are decreasing functions of $\mu_{0}(a\mid b)$ and $\mu_{0}(b)$ respectively, So if $b^{\prime}$ is the minimum probability symbol and $(b^{\prime},a^{\prime})$ is the minimum probability symbol pair then we can bound the above expression for $f_{n}(x_{n_{0}+1}^{n_{0}+n})$ as (for $\delta\leq{\mu_{0}(a\mid b)}$ and $\delta^{\prime}\leq{\mu_{0}(b)}$):
$f_{n}(x_{n_{0}+1}^{n_{0}+n})\leq\log\left(\frac{\mu_{0}(a^{\prime}\mid b^{\prime})}{\mu_{0}(a^{\prime}\mid b^{\prime})-\delta}\right)+\log(\frac{\mu_{0}(b^{\prime})}{\mu_{0}(b^{\prime})-\delta^{\prime}})$
$f_{n}(x_{n_{0}+1}^{n_{0}+n})\leq\log(\beta^{\prime})+\log(\beta^{\prime\prime})=\log(\beta)$
Lemma 3:The probability of error (false alarm rate) for the proposed test is bounded above and given by
$\mu_{0}(N(\alpha)<\infty)\leq\frac{\alpha}{2^{\lambda-\log\beta}-1}+\epsilon_{0}$
Where $\epsilon_{0}\to 0$ and $\beta\to 1$ as $n_{0}\to\infty$.
@@
$\leq\sum_{x_{1}^{n_{0}+n}\in B_{n}\cap E_{n}}(\hat{\mu}_{0}\!\left(x_{n_{0}+1}^{n_{0}+n}\right)\cdot 2^{\,n\log\beta}\bigr)$
$\leq\sum_{x_{1}^{n_{0}+n}\in B_{n}\cap E_{n}}2^{\log\alpha-L(x_{n_{0}+1}^{n_{0}+n})\,-n(\lambda-\log\beta)}$
$\alpha\cdot 2^{\,-n(\lambda-\log\beta)}(\sum_{x_{1}^{n_{0}+n}\in B_{n}\cap E_{n}}2^{-L(x_{n_{0}+1}^{n_{0}+n})})$
$\leq\alpha\cdot 2^{\,-n(\lambda-\log\beta)}$
The second inequality comes from definition of $f_{n}(x_{n_{0}+1}^{n_{0}+n})$ while the third inequality is due to Lemma 2 as $x_{1}^{n_{0}+n}\in E_{n}$.The fourth inequality comes from definition of $B_{n}$ while the last inequality comes from Kraft’s inequality. As $P(A\cap B)\leq{P(B)}$,we have
$\mu_{0}(B\cap(E)^{c})\leq{\mu_{0}((E)^{c})}\leq{\epsilon_{0}}$
Combining both equations, we have
$\mu_{0}(N(\alpha)<\infty)\leq{\alpha}\sum_{n=1}^{\infty}2^{-n(\lambda-\log\beta)}+\epsilon_{0}$
$\mu_{0}(N(\alpha)<\infty)\leq\frac{\alpha}{2^{\lambda-\log\beta}-1}+\epsilon_{0}$
By comparing the probability of error expressions in Lemma 3 and for actual pre-change case (JB-Page CUSUM Test), we observe that the probability of error increases in the case of the empirical pre-change distribution test. The probability of error now becomes a function of $\log\beta$. For the probability of error to be bounded, we require:
$\lambda>\max(\log\beta,D(\mu_{0}||{\widehat{\mu}}_{0}))$
In contrast, for the actual pre-change distribution case, $\lambda>0$ would have sufficed. However, as $n_{0}\to\infty$, the lower bound $\max(\log\beta,D(\mu_{0}||{\widehat{\mu}}_{0}))$ and the error term $\epsilon_{0}$ can be made arbitrarily small.
@@
$\max(\log\beta,D(\mu_{0}||{\widehat{\mu}}_{0}))<\lambda<D(\mu_{1}||\hat{\mu}_{0})$
@@ III-D Behaviour Under Postchange Distribution
Lemma 4: Let the conditions of Lemma 1 hold true.The proposed test will terminate with probability 1 under $\mu_{1}$ if $\lambda<D(\mu_{1}||{\widehat{\mu}}_{0})$ i.e.,
@@
Theorem 2: Let $\mu_{0}$ and $\mu_{1}$ be stationary and ergodic finite order Markov sources satisfying (2.10),(2.16),(2.17) and $\mu_{1}\ll\mu_{0}$. For the Modified CUSUM Test (3.7) with codes satisfying (2.7) and stopping time $N(\alpha)$ (3.10) as $\alpha\to 0$, for $\lambda<D(\mu_{1}||{\widehat{\mu}}_{0})$, we have
@@ IV Performance Bounds
${E_{0}}\left(M(\gamma)\right)\geq\frac{1}{2}+\frac{\gamma}{\frac{2}{2^{\lambda-\log\beta}-1}+2\epsilon_{0}\gamma}>\frac{\gamma}{\frac{2}{2^{\lambda-\log\beta}-1}+2\epsilon_{0}\gamma}$
@@
Proof: Equation (4.9) follows from (3.15) of Lemma 3 and (4.5) of Theorem 3(restatement of Theorem 1 from [16]) Equation (4.10) follows from (4.6) of Theorem 3(restatement of Theorem 1 from [16]) and Theorem 2.
@@
$\eta=\gamma\left(\frac{1}{2^{\lambda-\log(\beta)}-1}+\epsilon_{0}\gamma\right).$
@@
$\eta=\gamma\left(\frac{1}{2^{\lambda-\log(\beta)}-1}+1\right).$
@@
$\overline{E_{1}}(M\mid x_{1}^{n_{0}})\sim\frac{\log\gamma}{D(\mu_{1}||\widehat{\mu_{0}})-\lambda}$
@@
Hence, as $n_{0}\to\infty$, we get

### v2

@@ II-A Previous Work
Here, $\lambda$ was introduced by the authors in order to ensure a negative drift in the test statistic in the pre-change phase, as is the case in the classical Page’s test in the parametric setup.
@@ III-A Estimating Pre-change Distribution
$\hat{\mu}_{0}(x_{1}^{n_{0}})=\frac{\hat{\mu}_{0}(x_{1},x_{2})\hat{\mu}_{0}(x_{2},x_{3})\dots\hat{\mu}_{0}(x_{n_{0}-1},x_{n_{0}})}{\hat{\mu}_{0}(x_{2})\dots\hat{\mu}_{0}(x_{n_{0}-1})}$
@@
Here, $N(b)\mid x_{1}^{n_{0}}$ represents the number of times the symbol $b$ appears in the sequence $x_{1}^{n_{0}}$, whereas $N(b,a)\mid x_{1}^{n_{0}}$ represents the number of times the symbols $b,a$ jointly appear in the sequence $x_{1}^{n_{0}}$.Also, $\hat{\mu}_{0}(b)$ is the fraction of times $b$ appears in the sequence $x_{1}^{n_{0}}$, whereas $\hat{\mu}_{0}(b,a)$ is the fraction of times $(b,a)$ appear in the sequence $x_{1}^{n_{0}}$ jointly.
We assume that $n_{0}$ is large enough so that each symbol $b\in\mathcal{X}$ which has non-zero probability under ${\mu}_{0}$ has appeared at least once in $x_{1}^{n_{0}}$ and each symbol pair $(b,a)\in\mathcal{X}^{2}$ which has non-zero probability under ${\mu}_{0}$ has appeared at least once in $x_{1}^{n_{0}}$ (else $\hat{\mu}_{0}$ is not an accurate estimate of ${\mu}_{0}$).
Since ${\mu}_{0}$ is a stationary and ergodic finite-order Markov source, by Ergodic Theorem, we can say that there exists an $n_{0}$ such that,
@@ III-C Behaviour Under Prechange Distribution
Lemma 2: If $x_{1}^{n_{0}+n}\in E_{n}$, then $f_{n}(x_{n_{0}+1}^{n_{0}+n})$ is bounded above for all $n$ as $f_{n}\leq W_{n}$ where $\beta_{1}=\frac{\mu_{0}(a^{\prime}\mid b^{\prime})}{\mu_{0}(a^{\prime}\mid b^{\prime})-\delta},\beta_{2}=\frac{\mu_{0}(b^{\prime})}{\mu_{0}(b^{\prime})-\delta^{\prime}}$ and $(b^{\prime},a^{\prime})$ is such that $(b^{\prime},a^{\prime})$ is the minimum transition probability symbol pair and $b^{\prime}$ is such that $b^{\prime}$ is the minimum probability symbol.As $\delta\to 0$$\delta^{\prime}\to 0$ and $n_{0}\to\infty$, we have $\beta_{1}\to 1$ and $\beta_{2}\to 1$ .
@@
The terms on RHS are decreasing functions of $\mu_{0}(a\mid b)$ and $\mu_{0}(b)$ respectively, So if $b^{\prime}$ is the minimum probability symbol and $(b^{\prime},a^{\prime})$ is the minimum transition probability symbol pair then we can bound the above expression for $f_{n}(x_{n_{0}+1}^{n_{0}+n})$ as (for $\delta\leq{\mu_{0}(a\mid b)}$ and $\delta^{\prime}\leq{\mu_{0}(b)}$):
$f_{n}(x_{n_{0}+1}^{n_{0}+n})\leq\frac{n-1}{n}\cdot\log\left(\frac{\mu_{0}(a^{\prime}\mid b^{\prime})}{\mu_{0}(a^{\prime}\mid b^{\prime})-\delta}\right)+\frac{1}{n}\cdot\log(\frac{\mu_{0}(b^{\prime})}{\mu_{0}(b^{\prime})-\delta^{\prime}})$
$f_{n}(x_{n_{0}+1}^{n_{0}+n})\leq\frac{(n-1)\log(\beta_{1})+\log(\beta_{2})}{n}=W_{n}$
Lemma 3:The probability of error (false alarm rate) for the proposed test is bounded above by,
$\mu_{0}(N(\alpha)<\infty)\leq\frac{\alpha}{\dfrac{\beta_{1}\left(2^{\lambda-\log\beta_{1}}-1\right)}{\beta_{2}}}+\epsilon_{0}$
where $\epsilon_{0}\to 0$$\beta_{1}\to 1$ and $\beta_{2}\to 1$ as $n_{0}\to\infty$.
@@
$\leq\sum_{x_{1}^{n_{0}+n}\in B_{n}\cap E_{n}}(\hat{\mu}_{0}\!\left(x_{n_{0}+1}^{n_{0}+n}\right)\cdot 2^{\,nW_{n}}\bigr)$
$\leq\sum_{x_{1}^{n_{0}+n}\in B_{n}\cap E_{n}}2^{\log\alpha-L(x_{n_{0}+1}^{n_{0}+n})\,-n(\lambda-W_{n})}$
$\alpha\cdot 2^{\,-n(\lambda-W_{n})}(\sum_{x_{1}^{n_{0}+n}\in B_{n}\cap E_{n}}2^{-L(x_{n_{0}+1}^{n_{0}+n})})$
$\leq\alpha\cdot 2^{\,-n(\lambda-W_{n})}$
(3.20) follows from (3.16) while the third inequality is due to Lemma 2 as $x_{1}^{n_{0}+n}\in E_{n}$.The fourth inequality comes from definition of $B_{n}$ while the last inequality comes from Kraft’s inequality. As $P(A\cap B)\leq{P(B)}$,and from (3.5) and (3.6), we have,
$\mu_{0}(B\cap(E)^{c})\leq{\mu_{0}((E)^{c})}\leq{\epsilon}+{\epsilon^{\prime}}={\epsilon_{0}}$
Combining both equations and from (3.14) we have,
$\mu_{0}(N(\alpha)<\infty)\leq{\alpha}\sum_{n=1}^{\infty}2^{-n(\lambda-\log\beta_{1})+\log(\frac{\beta_{2}}{\beta_{1}})}+\epsilon_{0}$
$\mu_{0}(N(\alpha)<\infty)\leq\frac{\alpha}{\dfrac{\beta_{1}\left(2^{\lambda-\log\beta_{1}}-1\right)}{\beta_{2}}}+\epsilon_{0}$
By comparing the probability of error expressions in Lemma 3 and for actual pre-change case (2.19), we observe that the probability of error increases in the case of the empirical pre-change distribution test. The probability of error now becomes a function of $\log\beta_{1}$$\beta_{1}$ and $\beta_{2}$. For the probability of error to be bounded, we require:
$\lambda>\max(\log\beta_{1},D(\mu_{0}||{\widehat{\mu}}_{0}))$
In contrast, for the actual pre-change distribution case, $\lambda>0$ would have sufficed. However, as $n_{0}\to\infty$, the lower bound $\max(\log\beta_{1},D(\mu_{0}||{\widehat{\mu}}_{0}))$,the error term $\epsilon_{0}$ and $\beta_{1},\beta_{2}$ can be made as small as needed.
@@
$\max(\log\beta_{1},D(\mu_{0}||{\widehat{\mu}}_{0}))<\lambda<D(\mu_{1}||\hat{\mu}_{0})$
@@ III-D Behaviour Under Postchange Distribution
Lemma 4: Let the conditions of Lemma 1 hold true.The proposed test (3.7) will terminate with probability 1 under $\mu_{1}$ if $\lambda<D(\mu_{1}||{\widehat{\mu}}_{0})$ i.e.,
@@
Theorem 2: Let $\mu_{0}$ and $\mu_{1}$ be stationary and ergodic finite order Markov sources satisfying (2.10),(2.16),(2.17) and $\mu_{1}\ll\mu_{0}$. For the Modified CUSUM Test (3.7) with codes satisfying (2.7) and stopping time $N(\alpha)$ (3.10) with $\alpha\to 0$, for $\lambda<D(\mu_{1}||{\widehat{\mu}}_{0})$, we have
@@ IV Performance Bounds
${E_{0}}\left(M(\gamma)\right)\geq\frac{1}{2}+\frac{\gamma}{\frac{2\beta_{2}}{\beta_{1}(2^{\lambda-\log\beta_{1}}-1)}+2\epsilon_{0}\gamma}>\frac{\gamma}{\frac{2\beta_{2}}{\beta_{1}(2^{\lambda-\log\beta_{1}}-1)}+2\epsilon_{0}\gamma}$
@@
Proof: Equation (4.9) follows from (3.15) of Lemma 3 and (4.5) of Theorem 3(restatement of Theorem 1 from [16]). Equation (4.10) follows from (4.6) of Theorem 3(restatement of Theorem 1 from [16]) and Theorem 2.
@@
$\eta=\gamma\left(\frac{\beta_{2}}{\beta_{1}(2^{\lambda-\log(\beta_{1})}-1)}+\epsilon_{0}\gamma\right).$
@@
$\eta=\gamma\left(\frac{\beta_{2}}{\beta_{1}(2^{\lambda-\log(\beta_{1})}-1)}+1\right).$
@@
$\overline{E_{1}}(M\mid x_{1}^{n_{0}})\sim\frac{\log\eta}{D(\mu_{1}||\widehat{\mu_{0}})-\lambda}$
@@
Hence, as $n_{0}\to\infty$ and because samples $x_{1}^{n_{0}}$ are generated using $\mu_{0}$, we get